# Integrating Weather Station Data and Radar for Precipitation Nowcasting: SmaAt-fUsion and SmaAt-Krige-GNet

### v1

@@ Abstract
In recent years, data-driven, deep learning-based approaches for precipitation nowcasting have attracted significant attention, showing promising results. However, many existing models fail to fully exploit the extensive atmospheric information available, relying primarily on precipitation data alone. This study introduces two novel deep learning architectures, SmaAt-fUsion and SmaAt-Krige-GNet, specifically designed to enhance precipitation nowcasting by integrating multi-variable weather station data with radar datasets. By leveraging additional meteorological information, these models improve representation learning in the latent space, resulting in enhanced nowcasting performance. The SmaAt-fUsion model extends the SmaAt-UNet framework by incorporating weather station data through a convolutional layer, integrating it into the bottleneck of the network. Conversely, the SmaAt-Krige-GNet model combines precipitation maps with weather station data processed using Kriging, a geo-statistical interpolation method, to generate variable-specific maps. These maps are then utilized in a dual-encoder architecture based on SmaAt-GNet, allowing multi-level data integration. Experimental evaluations were conducted using four years (2016–2019) of weather station and precipitation radar data from the Netherlands. Results demonstrate that SmaAt-Krige-GNet outperforms the standard SmaAt-UNet, which relies solely on precipitation radar data, in low precipitation scenarios, while SmaAt-fUsion surpasses SmaAt-UNet in both low and high precipitation scenarios. This highlights the potential of incorporating discrete weather station data to enhance the performance of deep learning-based weather nowcasting models.
@@ 1 Introduction
Precipitation nowcasting refers to the short-term forecasting of rainfall intensity, typically within a timeframe ranging from a few minutes to several hours. Accurate nowcasting is crucial for various weather-dependent sectors, including flood management, transportation, energy distribution, and emergency response [1, 2]. The effectiveness of nowcasting models depends on their ability to capture the complex, dynamic nature of precipitation patterns across different spatial and temporal scales.
Traditional precipitation nowcasting approaches rely on numerical weather prediction (NWP) methods. NWP models utilize physical atmospheric properties to generate multiple precipitation scenarios, but they are computationally expensive and often impractical for short-term forecasts due to their sensitivity to initial conditions and require high computations [3].
@@
Recent advancements in deep learning have led to the development of advanced data-driven models that utilize historical weather data to capture complex spatiotemporal patterns [4, 5, 6]. Convolutional Neural Networks (CNNs) based models, in particular, have demonstrated promising results in precipitation nowcasting by effectively extracting spatial features from radar images [7, 8]. Despite their success, many existing deep learning models for precipitation nowcasting primarily rely on radar-based precipitation data, often overlooking the potential benefits of integrating additional meteorological information from other sources, such as weather stations.
@@
Some recent studies have attempted to incorporate multiple atmospheric variables into their models. For instance, the authors in [9] introduced WF-UNet, a two-stream UNet architecture that processes both precipitation maps and wind speed maps to assess the impact of wind on precipitation nowcasting. While this approach leverages multi-variable meteorological data, it relies entirely on spatially continuous maps rather than integrating discrete station-based observations. In contrast, our work aims to bridge this gap by learning from both discrete weather station data and radar-based precipitation maps. Weather stations provide highly localized, multi-variable atmospheric measurements (e.g., temperature, humidity, wind speed, and pressure), offering valuable insights into weather dynamics. However, station data alone lacks spatial continuity, making it challenging to use directly in deep learning models designed for grid-based data.
@@
To address this limitation, this study explores the potential benefits of incorporating additional meteorological data into deep learning-based nowcasting models. Specifically, we introduce two novel architectures, i.e. SmaAt-fUsion and SmaAt-Krige-GNet, that integrate multi-variable weather station data alongside precipitation radar data. The key contributions of this work are as follows:
@@
    Introduction of a novel dataset: We have collected a comprehensive dataset containing atmospheric data from 2016-2019 for 22 weather stations in the Netherlands, each recording 8 meteorological variables. This dataset provides a rich source of information for improving precipitation nowcasting.
@@
    Development of SmaAt-fUsion: This model extends the SmaAt-UNet framework [10] by incorporating weather station data into the bottleneck of the SmaAt-UNet model, where high-level representations are formed.
@@ 2 Related Work
Recent advancements in data-driven weather forecasting have increasingly leveraged deep learning techniques to improve weather elements forecasting and nowcasting [11, 12, 13, 14, 15]. Various neural network architectures have been successfully applied in weather forecasting, including Recurrent Neural Networks (RNNs) [16], Long Short-Term Memory (LSTM) networks [17], Convolutional LSTM (ConvLSTM) [18], Convolutional Neural Networks (CNNs) [8], Graph Neural Networks (GNNs) [19], and encoder-decoder frameworks [20].
@@
More recently, the UNet architecture [21], originally developed for medical image segmentation, has demonstrated effectiveness in precipitation due to its encoder-decoder structure with skip connections, enabling the preservation of fine-grained spatial details. Variations of UNet have been proposed to enhance performance while maintaining computational efficiency for weather prediction. For example, SmaAt-UNet [10], an enhanced version of the UNet model that integrates attention modules and depthwise-separable convolutions, substantially decreases the number of trainable parameters in the traditional UNet while maintaining its performance. SelfAttUNet [22] integrates a self-attention mechanism into the UNet model, allowing it to highlight key regions in an image for more effective radar-based precipitation nowcasting. Several other studies, such as [9, 23, 24], have also explored improvements to deep learning models built on the core UNet architecture, highlighting their effectiveness in weather nowcasting tasks.
@@
In addition to CNN-based architectures, data fusion techniques have been increasingly explored to integrate multiple meteorological data sources for enhanced precipitation nowcasting. Traditional models primarily rely on radar reflectivity data; however, recent studies have demonstrated that incorporating additional atmospheric variables, such as wind speed, temperature, and humidity, significantly improves short-term forecasting accuracy. By leveraging multiple data sources, fusion models better capture complex weather dynamics, leading to more reliable nowcasting performance. Kaparakis et al. [9] proposed the Weather Fusion UNet (WF-UNet), a fusion-based deep learning model for precipitation nowcasting over Western Europe. Their findings underscore the role of fusion models in enhancing short-term precipitation predictions by leveraging complementary data sources. Similarly, Leinonen et al. [25] introduced a deep learning-based thunderstorm nowcasting model that integrates diverse meteorological data, including weather radar, lightning detection, satellite imagery, numerical weather prediction outputs, and digital elevation models. Furthermore, the Spatiotemporal Feature Fusion Transformer [26] incorporates feature-level fusion mechanisms to enhance precipitation nowcasting performance. Collectively, these works illustrate the growing impact of data fusion techniques in weather nowcasting, highlighting their potential to bridge the gap between numerical weather prediction and deep learning-based methodologies.
@@ 3.1 Dataset
This study utilizes two complementary datasets: a radar-based precipitation dataset and a weather station dataset. Both datasets were obtained from the Royal Netherlands Meteorological Institute (KNMI 111 https://www.knmi.nl/home.) and are used together to improve precipitation nowcasting performance.
Precipitation Dataset: The precipitation dataset, originally used by Trebing et al. [10], consists of radar-based rainfall data collected over four years (2016–2019) in the Netherlands. It contains approximately 420,000 rain maps, each captured at 5-minute intervals. The data is sourced from two C-band Doppler weather radar stations located in: De Bilt (52.10°N, 5.18°E, 44 m MSL) and Den Helder (52.96°N, 4.79°E, 51 m MSL). To generate rain maps, each radar performs four full 360° azimuthal scans at beam elevation angles of 0.3°, 1.1°, 2.0°, and 3.0° [27, 10].
Weather Station Dataset: The second dataset consists of weather station data, which provides multivariable atmospheric measurements from 22 weather stations across the Netherlands (see Fig. 1). This dataset is novel and contains 8 meteorological variables, selected based on availability, relevance, and consistency. The recorded variables are: Temperature, Humidity, Atmospheric Pressure, Wind Speed, Max Wind Speed, Wind Speed Deviation, Wind Direction, Wind Direction Deviation.
@@
Dataset Integration and Processing: To create a unified dataset for model input, the precipitation and weather station datasets undergo several preprocessing steps to ensure consistency and alignment. For the precipitation maps, a 288×288 pixel cutout is taken from the center of the Netherlands, covering the area with the highest data availability. The selected geographic boundaries are: Longitude: 4.29°E to 6.73°E and Latitude: 51.32°N to 52.81°N. These maps are then resized to 64×64 pixels to reduce memory requirements, as both datasets are incorporated into the models. Finally, the precipitation values are normalized using the maximum precipitation value in the dataset (47.83 mm). The precipitation dataset contains a significant number of radar images with little to no recorded rainfall. To mitigate the risk of biasing the model toward predicting zero values, precipitation maps where fewer than 50% of the pixels indicate rainfall are excluded. Once filtered, the remaining precipitation maps are temporally matched with the corresponding weather station data recorded at the same timestamp. Any data point that does not have a corresponding entry in the other dataset is removed, ensuring that only fully aligned data pairs are retained. To maintain consistency across variables, all weather station measurements are standardized based on their respective means and standard deviations. A challenge in merging the two datasets arises from their differing temporal resolutions. While precipitation maps are recorded at 5-minute intervals, weather station data is available at 10-minute intervals. To resolve this mismatch, the weather station dataset is duplicated to align with the precipitation timestamps, allowing for seamless integration of both sources. After preprocessing, the final dataset is divided into a training set (2016–2018) and a testing set (2019). During model training, a validation set is created by randomly selecting 10% of the training data.
Figure 2: Overview of the SmaAt-fUsion architecture. Weather station data is integrated into the model by concatenating its learned representation with that of the precipitation maps at the bottleneck of the SmaAt-UNet model.
@@
Rain Threshold Train Size Test Size Subset
$0\%$ (Original) 314940 105003 100%
$50\%$ 5734 1557 1.74%
@@ 3.2 Proposed models
In the following, we introduce two architectures designed to extend the core SmaAt-UNet model [10] to learn from both weather station data and precipitation maps. The key difference between these models lies in how they integrate weather station data into SmaAt-UNet, which primarily operates on gridded precipitation maps. Both models take 12 precipitation maps as input, representing one hour of weather radar data. In addition to these maps, they incorporate station data using different integration strategies. The output of each model is a single precipitation map predicting conditions 30 minutes after the last input frame.
@@ 3.3 SmaAt-fUsion model
In SmaAt-fUsion, our aim is to integrate weather station data into the core SmaAt-Unet model [10], which operates on gridded precipitation radar maps. To this end, following the approach of [4], we exploit the spatio-temporal structure of station data by transforming each regressor vector, containing historical weather variable measurements from all stations, into a tensor with dimensions (stations = 22, variables = 8, lags = 12). This tensor is processed through a double 3D depthwise separable convolution block, consisting of a depthwise convolution followed by a pointwise convolution. Compared to classical 3D convolutions, this decomposition significantly reduces the number of parameters while preserving spatial and temporal correlations in the data. The output from this convolutional block is then passed through a 3D Adaptive Max Pooling layer, which reshapes it to match the bottleneck tensor size of the original SmaAt-Unet model. Finally, the pooled representation is concatenated with the bottleneck tensor, as illustrated in Fig. 2. By integrating station data, we enrich the model’s latent space with additional insights from weather station observations. This fusion strengthens the model’s ability to better capture complex spatio-temporal dependencies, ultimately improving its performance in precipitation mapping.

### v2

@@ Abstract
Short-term precipitation nowcasting is essential for flood management, transportation, energy system operations, and emergency response. However, many existing models fail to fully exploit the extensive atmospheric information available, relying primarily on precipitation data alone. This study examines whether integrating multi-variable weather-station measurements with radar can enhance nowcasting skill and introduces two complementary architectures that integrate multi-variable station data with radar images. The SmaAt-fUsion model extends the SmaAt-UNet framework by incorporating weather station data through a convolutional layer, integrating it into the bottleneck of the network; The SmaAt-Krige-GNet model combines precipitation maps with weather station data processed using Kriging, a geo-statistical interpolation method, to generate variable-specific maps. These maps are then utilized in a dual-encoder architecture based on SmaAt-GNet, allowing multi-level data integration . Experimental evaluations were conducted using four years (2016–2019) of weather station and precipitation radar data from the Netherlands. Results demonstrate that SmaAt-Krige-GNet outperforms the standard SmaAt-UNet, which relies solely on precipitation radar data, in low precipitation scenarios, while SmaAt-fUsion surpasses SmaAt-UNet in both low and high precipitation scenarios. This highlights the potential of incorporating discrete weather station data to enhance the performance of deep learning-based weather nowcasting models.
@@ 1 Introduction
Precipitation nowcasting refers to the short-term forecasting of rainfall intensity, typically within a timeframe ranging from a few minutes to several hours. Accurate nowcasting is crucial for various weather-dependent sectors, including flood management, transportation, energy distribution, and emergency response li2024quantitative; wilson2010nowcasting. However, although numerical weather prediction (NWP) and radar-based methods have made notable progress, there remains improvement in operational settings. NWP models utilize physical atmospheric properties to generate multiple precipitation scenarios, but they are computationally expensive and often impractical for short-term forecasts due to their sensitivity to initial conditions and require high computations bauer2015quiet; sun2014use; wilks2011statistical. This gap poses a growing threat to the effectiveness of early warning systems, particularly in densely populated or infrastructure-critical regions such as the Netherlands. The effectiveness of nowcasting models depends on their ability to capture the complex, dynamic nature of precipitation patterns across different spatial and temporal scales.
@@
Recent advancements in deep learning have led to the development of advanced data-driven models that utilize historical weather data to capture complex spatiotemporal patterns mehrkanoon2019deep; ren2021deep; vatamany2025graph; price2025probabilistic; ham2023anthropogenic; allen2025end. Convolutional Neural Networks (CNNs) based models, in particular, have demonstrated promising results in precipitation nowcasting by effectively extracting spatial features from radar images ayzel2019all; ayzel2020rainnet; jiang2024hybrid. However, these models almost exclusively rely on radar reflectivity data, overlooking other atmospheric drivers, such as temperature, humidity, pressure, and wind, which fundamentally influence precipitation formation. As a result, they are unable to capture the localized signals that often precede rainfall initiation. Consequently, radar-only systems may fail to capture pre-convective signals, limiting both accuracy and generalization under rapidly changing weather conditions.
@@
Some recent studies have attempted to incorporate multiple atmospheric variables into their models to improve prediction robustness wang2025enhancing; cao2025enhancing. For instance, the authors in kaparakis2023wf introduced WF-UNet, a two-stream UNet architecture that processes both precipitation maps and wind speed maps to assess the impact of wind on precipitation nowcasting. However, such approaches typically depend on spatially continuous maps rather than integrating discrete station-based observations. Weather stations provide highly localized, multi-variable atmospheric observations that offer valuable insights into the physical processes behind rainfall. However, station data alone lacks spatial continuity, making it challenging to use directly in deep learning models designed for grid-based data. Our work aims to bridge this gap by learning from both discrete weather station data and radar-based precipitation maps.
Among various spatial interpolation techniques, Kriging Isaaks1989 has been widely used as a geostatistical method for generating spatially continuous fields from discrete observations. It estimates both spatial values and associated uncertainties through a variogram model 8102132; brown2022variogram. This makes Kriging performs well for smoothly varying meteorological variables and provides interpretable spatial correlations. However, it also suffers from several weaknesses when applied to high-resolution or rapidly evolving rainfall processes. It assumes spatial stationarity, leading to performance degradation during convective events; it tends to over-smooth local extremes, thereby reducing the representation of sharp rainfall gradients. These strengths and weaknesses highlight the need to combine Kriging interpolation with the deep learning to produce more realistic and dynamically consistent precipitation fields.
@@
To address this limitation, this study explores the potential benefits of incorporating additional meteorological data into deep learning-based nowcasting models. Specifically, we introduce two novel architectures, i.e. SmaAt-fUsion and SmaAt-Krige-GNet, that integrate multi-variable weather station data alongside precipitation radar data. The key contributions of this work are as follows.
@@
    Integrate multi-source dataset: We integrate four years (2016–2019) of radar-based precipitation data with multi-variable observations from 22 KNMI weather stations across the Netherlands. The datasets were temporally aligned and standardized to enable pixel-level fusion between radar grids and discrete station data, provides a rich source of information for improving precipitation nowcasting.
@@
    Development of SmaAt-fUsion: This model extends the SmaAt-UNet framework trebing2021smaat by incorporating weather station data into the bottleneck of the SmaAt-UNet model, where high-level representations are formed.
@@ 2 Related Work
Quantitative Precipitation Nowcasting (QPN) has experienced a fundamental transformation from statistical extrapolation and optical-flow approaches toward deep learning-based forecasting 10220887. Classical optical-flow systems such as RainyMotion ayzel2019optical and PySTEPS pulkkinen2019pysteps estimate the displacement of radar echoes between consecutive frames and have long served as operational baselines due to their simplicity and physical interpretability. Early deep learning architectures have been successfully applied in weather forecasting, such as Recurrent Neural Networks (RNNs) singh2019weather, Long Short-Term Memory (LSTM) networks zaytar2016sequence, Convolutional LSTM (ConvLSTM) shi2015convolutional, Convolutional Neural Networks (CNNs) ayzel2020rainnet, Graph Neural Networks (GNNs) simeunovic2021spatio, and encoder-decoder frameworks larraondo2019data. All of which have shown promising results in capturing complex spatio-temporal precipitation patterns.
@@
Building upon these advances, the UNet architecture ronneberger2015unetconvolutionalnetworksbiomedical, originally developed for medical image segmentation, has demonstrated remarkable effectiveness in precipitation nowcasting due to its encoder–decoder structure with skip connections, which enables the preservation of fine-grained spatial details. Several variations of UNet have been proposed to improve both performance and computational efficiency for weather prediction tasks. For instance, SmaAt-UNet trebing2021smaat, an enhanced lightweight version of UNet integrating attention modules and depthwise-separable convolutions, substantially reduces the number of trainable parameters while maintaining high predictive accuracy. SelfAttUNet nie2021self incorporates a self-attention mechanism to highlight key spatial regions for more effective radar-based precipitation prediction. Other studies, such as kaparakis2023wf; reulen2024ga; renault2023sar, have further extended the UNet architecture by combining attention, residual, and generative components, consistently confirming the suitability of encoder–decoder architectures for data-driven nowcasting.
@@
In addition to CNN-based architectures, multi-source data fusion techniques have been increasingly explored to integrate multiple meteorological data sources for enhanced precipitation nowcasting. Traditional models primarily rely on radar reflectivity data; however, recent studies have demonstrated that incorporating additional atmospheric variables, such as wind speed, temperature, and humidity, significantly improves short-term forecasting accuracy. Using multiple data sources, fusion models better capture complex weather dynamics, leading to more reliable nowcast performance. Kaparakis et al. kaparakis2023wf proposed the Weather Fusion UNet (WF-UNet) integrates radar reflectivity and wind maps through a dual-stream UNet framework, demonstrating that multi-variable fusion enhances nowcasting accuracy. leinonen2023thunderstorm introduced a deep learning-based thunderstorm nowcasting model that integrates diverse meteorological data, including weather radar, lightning detection, satellite imagery, numerical weather prediction outputs, and digital elevation models. Large-scale models such as MM-RNN ma2023mm and MetNet-3 andrychowicz2023deep follow this trend by combining radar, satellite, and numerical weather prediction (NWP) data in a single framework to learn relationships across different spatial and temporal scales and to produce more physically consistent forecasts. Furthermore, the Spatiotemporal Feature Fusion Transformer xiong2024spatiotemporal incorporates feature-level fusion mechanisms to enhance precipitation nowcasting performance. Collectively, these works illustrate the growing impact of data fusion techniques in weather nowcasting, highlighting their potential to bridge the gap between numerical weather prediction and deep learning-based methodologies.
@@ 3.1 Dataset
This study utilizes two complementary datasets: a radar-based precipitation dataset and a weather station dataset. Both datasets were obtained from the Royal Netherlands Meteorological Institute (KNMI  [^3] https://www.knmi.nl/home.) and are used together to improve precipitation nowcasting performance.
Precipitation Dataset: The precipitation dataset, originally used by Trebing et al. trebing2021smaat, consists of radar-based rainfall data collected over four years (2016–2019) in the Netherlands. It contains approximately 420,000 rain maps, each captured at 5-minute intervals. The data is sourced from two C-band Doppler weather radar stations located in: De Bilt (52.10°N, 5.18°E, 44 m MSL) and Den Helder (52.96°N, 4.79°E, 51 m MSL). The technical characteristics of these radars including a wavelength of Wavelength (cm), Pulse repetition frequency (Hz), Peak power (kW), Pulse duration ($\mu s$), 3-db beamwidth, Antenna rotation speed of $18\textdegree s^{-1}$, and Number of samples per range ($km^{-1}$). These specifications determine the spatial resolution, sensitivity, and data quality of the radar observations. To generate rain maps, each radar performs four full 360° azimuthal scans at beam elevation angles of 0.3°, 1.1°, 2.0°, and 3.0° precipitation_dataset; trebing2021smaat. The rain maps are rain-gauge adjusted with more details being described in overeem2009derivation. This is the same radar-derived precipitation dataset used in Trebing et al. (2021) and serves as the precipitation input for the models in this study.
Weather Station Dataset: The second dataset consists of weather station data, which provides multivariable atmospheric measurements from 22 weather stations across the Netherlands (see Fig. 1).This dataset contains 8 meteorological variables, selected based on availability, relevance, and consistency. The recorded variables are: Air Temperature, Humidity, Atmospheric Pressure, Average wind Speed, Wind Direction, Standard Deviation Wind Direction, Standard Deviation Wind Direction, Max Wind Speed.
@@
Dataset Integration and Processing: To create a unified dataset for model input, the precipitation and weather station datasets undergo several preprocessing steps to ensure consistency and alignment. The eight meteorological variables recorded at each station capture local atmospheric conditions that influence the formation and movement of precipitation. These variables provide complementary information that is not present in the radar sequence alone. For the precipitation maps, a 288×288 pixel cutout is taken from the center of the Netherlands, covering the area with the highest data availability. The selected geographic boundaries are: Longitude: 4.29°E to 6.73°E and Latitude: 51.32°N to 52.81°N. These maps are then resized to 64×64 pixels to reduce memory requirements, as both datasets are incorporated into the models. Finally, the precipitation values are normalized using the maximum precipitation value in the dataset (47.83 mm). The precipitation dataset contains a significant number of radar images with little to no recorded rainfall.
To mitigate the risk of biasing the model toward predicting zero values, precipitation maps where fewer than 50% of the pixels indicate rainfall are excluded. As summarized in Table 1, applying this filtering criterion reduces the full 4-year dataset from 314,940 training samples and 105,003 test samples to 5,734 and 1,557 samples, respectively. Once filtered, the remaining precipitation maps are temporally matched with the corresponding weather station data recorded at the same timestamp. Any data point that does not have a corresponding entry in the other dataset is removed, ensuring that only fully aligned data pairs are retained. To maintain consistency across variables, all weather station measurements are standardized based on their respective means and standard deviations. A challenge in merging the two datasets arises from their differing temporal resolutions. While precipitation maps are recorded at 5-minute intervals, weather station data is available at 10-minute intervals. To resolve this mismatch, the weather station dataset is duplicated to align with the precipitation timestamps, allowing for seamless integration of both sources.
In SmaAt-fUsion, the time series of all meteorological variables across all stations is transformed into a compact 3D tensor, allowing the model to learn how local atmospheric conditions modulate the high-level features at the bottleneck of the SmaAt-UNet. This fusion enriches the latent precipitation representation with complementary physical signals that are not contained in the radar sequence alone. In SmaAt-Krige-GNet, each meteorological variable at each timestep is interpolated using Kriging to create spatially continuous 64 × 64 maps. This process propagates local station measurements across the spatial domain, enabling the model to learn how spatial gradients in temperature, humidity, wind, and pressure relate to evolving precipitation structures. After preprocessing, the final dataset is divided into a training set (2016–2018) and a testing set (2019). During model training, a validation set is created by randomly selecting 10% of the training data.
Figure 2: Overview of the SmaAt-fUsion architecture. Weather station data is integrated into the model by concatenating its learned representation with that of the precipitation maps at the bottleneck of the SmaAt-UNet model. The precipitation stream processes 12 radar maps through the standard SmaAt-UNet encoder–decoder: depthwise-separable convolutions (blue arrows), max-pooling (brown), bilinear upsampling (green), and skip connections (grey). Features progressively compress through the encoder and are then expanded during decoding, forming the main spatial–temporal pathway of the network. In parallel, the station-data stream transforms the 22×8×12 tensor using two 3D depthwise-separable convolutions (pink) and an AdaptiveMaxPool3D layer to produce a compact embedding. This embedding flows into the fusion stage, where it is concatenated with the UNet bottleneck, creating a shared latent representation used for the final precipitation prediction.
@@
Rain Threshold Train Size Test Size Subset $0\%$ (Original) 314940 105003 100% $50\%$ 5734 1557 1.74%
@@ 3.2 Proposed models
In the following, we introduce two architectures designed to extend the core SmaAt-UNet model trebing2021smaat to learn from both weather station data and precipitation maps. While both models take 12 radar images (representing one hour of weather radar data) as input and predict the precipitation field 30 minutes ahead, they differ in how the station observations are encoded and integrated. SmaAt-fUsion embeds the weather station variables directly into the UNet bottleneck, allowing meteorological variables to influence high-level latent features. SmaAt-Krige-GNet, in contrast, incorporate station measurements into spatially radar maps using Kriging and processes them through a parallel encoder, enabling multi-scale fusion across the entire UNet architecture.
@@ 3.3 SmaAt-fUsion model
In SmaAt-fUsion, our aim is to integrate weather station data into the core SmaAt-Unet model trebing2021smaat, which operates on gridded precipitation radar maps. To this end, following the approach of mehrkanoon2019deep, we exploit the spatio-temporal structure of station data by transforming each regressor vector, containing historical weather variable measurements from all stations, into a tensor with dimensions (stations = 22, variables = 8, lags = 12). This tensor is processed through a double 3D depthwise separable convolution block, consisting of a depthwise convolution followed by a pointwise convolution. Compared to classical 3D convolutions, this decomposition significantly reduces the number of parameters while preserving spatial and temporal correlations in the data. The output from this convolutional block is then passed through a 3D Adaptive Max Pooling layer, which reshapes it to match the bottleneck tensor size of the original SmaAt-Unet model. Finally, the pooled representation is concatenated with the bottleneck tensor, as illustrated in Fig. 2. By integrating station data, we enrich the model’s latent space with additional insights from weather station observations. This fusion strengthens the model’s ability to better capture complex spatio-temporal dependencies, ultimately improving its performance in precipitation mapping.